40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 24 21:02:46 2009, Alberto, Update, PSL, Aligning the beam to the Faraday 
    Reply  Sat Jul 25 16:04:32 2009, rob, Update, PSL, Aligning the beam to the Faraday 
       Reply  Sat Jul 25 19:04:01 2009, Koji, Update, PSL, Aligning the beam to the Faraday MC2_Cam.jpg
          Reply  Sun Jul 26 13:19:54 2009, rana, Update, PSL, Aligning the mode cleaner Untitled.png
             Reply  Sun Jul 26 16:05:17 2009, Alberto, Update, PSL, Aligning the mode cleaner 
             Reply  Mon Jul 27 09:34:07 2009, steve, Summary, IOO, Aligning the mode cleaner 20dtrend.jpg
Message ID: 1788     Entry time: Fri Jul 24 21:02:46 2009     Reply to this: 1791
Author: Alberto 
Type: Update 
Category: PSL 
Subject: Aligning the beam to the Faraday 

This afternoon I kept working on the alignment of the beam so that it matches at the same the PSL periscope, the Mode Cleaner and the Faraday isolator at the input of the IFO.

The camera looking at the Farady showed a beam quite low from the center of the Faraday's entrance. I wanted to move it up.

After working on the periscope alignment and on the MC mirrors, I think I managed to moved it up a bit. To know whether that was enough or not I wanted to evaluate the alignment to the X arm by checking the value of TRX.

In order for the MC to be finely matched to the input beam from the periscope, the WFS controls have to be on. Before turning them on, I centered the beam on their QPDs and run the WFS_zero_offset script.

When I turned them on, the control signal in Pitch from WFS2 started going up with no stop. It was like the integrator in the loop was fed with a DC bias. The effect of that was to misalign the MC cavity from the good state in which it was with the only length control on (that is, transmission ~2.7, reflection ~ 0.4).

I don't know why that is happening. To exclude that it was due to a computer problem I first burtrestored C1IOO to July the 18th, but since that did not help, I even restarted it. Also that didn't solve the problem.

 

Flashes at ETMX show at least that the beam is going through the Farady. How well, I can't tell untill the MC is under full control.

 

I have to leave the lab now, but I can be back tomorrow to keep working on that.

ELOG V3.1.3-